OpenStudy (anonymous):

Perform the indicated operations.y^4-16/y+2 divide y^2+4/5

OpenStudy (anonymous):

5(y - 4) 5(y + 4) 5(y - 2)

OpenStudy (anonymous):

Factor everything that you can so when you divide you can easily see if there are any factors that cancel one another out.

OpenStudy (anonymous):

i need help please

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@hero

OpenStudy (phi):

use (a^2 - b^2) = (a-b)(a+b) notice that y^4-16 matches this pattern. \(y^4 \) is \(y^2 \cdot y^2\) or \( (y^2)^2 \) and 16 is \(4^2\)

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@ saifoo.khan (Ambassador

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

\[\frac{y^4-16}{y+2}\div\frac{y^2+4}{5}\]\[\frac{y^4-16}{y+2}*\frac{5}{y^2+4}\]\[\frac{(y^2-4)(y^2+4)}{y+2}*\frac{5}{y^2+4}\]\[\text{So you get..}\]\[\frac{5(y^2-4)}{y+2}\]\[\text{Now solve and you would get the answer.}\]

OpenStudy (anonymous):

@robert2

OpenStudy (anonymous):

a;5(y - 4) b;5(y + 4) c;5(y - 2) what is tha anwrer i confusing

OpenStudy (anonymous):

\[\frac{5((y+2)(y-2)}{y+2}=5(y-2)\]

OpenStudy (anonymous):

is true

OpenStudy (anonymous):

If there is something confusing,ask me.. :)

OpenStudy (anonymous):

C is the answer..

OpenStudy (anonymous):

You are welcome..